OpenStudy (anonymous):

The formula S=n(a^1+a^n)/2 gives partial sum of an arithmetic sequence. What is the formula solved for a^n?

OpenStudy (anonymous):

i am going to make a guess that this says \[\large S=\frac{n(a_1+a_n)}{2}\] am i right?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

and you want to solve for \(a_n\)

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

multiply both side by 2 first and get \[2S=n(a_1+a_n)\]

OpenStudy (anonymous):

divide both sides by \(n\) and get \[\frac{2S}{n}=a_1+a_n\]

OpenStudy (anonymous):

then subtract \(a_1\) from both sides and you are done \[\large\frac{2S}{n}-a_1=a_n\]

OpenStudy (anonymous):

thank you